# LibEER: A Comprehensive Benchmark and Algorithm Library for EEG-based Emotion Recognition

### v2

@@ V Experimental Results
DBN MS-MDA HSLT EEGNet CDCN TSception ACRNN DGCNN RGNN GCBNet GCBNet_BLS
SEED ACC 37.07 36.16 64.00 56.00 38.19 57.72 45.60 45.39 60.87 57.20 56.32 56.32
F1 33.45 22.67 57.35 55.75 31.83 58.66 43.54 42.37 57.22 51.37 55.12 51.43
SEED-IV ACC 28.98 36.82 56.07 30.33 28.19 31.03 34.19 31.97 42.54 44.13 32.27 40.54
F1 24.56 32.60 48.68 11.64 28.35 27.01 26.83 18.82 43.10 43.30 32.89 42.73
HCI-V ACC 70.33 69.27 67.64 66.94 57.06 67.69 57.36 54.53 63.19 65.89 65.16 71.06
F1 66.36 65.51 53.70 62.22 53.83 62.67 54.76 52.58 58.75 44.33 63.32 61.83
HCI-A ACC 54.41 57.50 57.23 49.48 54.70 55.93 52.30 51.23 59.42 57.26 52.21 60.85
F1 52.91 56.24 55.68 47.43 54.02 55.15 50.25 49.45 57.02 56.89 51.84 56.26
HCI-VA ACC 31.56 28.30 44.92 35.07 34.84 30.09 26.99 27.21 41.54 37.43 43.02 36.92
F1 26.29 27.58 21.65 27.19 27.96 24.71 21.95 20.92 38.08 19.60 36.07 32.99
DEAP-V ACC 49.58 55.02 54.82 56.56 52.36 57.78 54.44 51.94 49.91 52.15 53.58 52.34
F1 48.49 53.14 50.60 56.56 49.74 57.72 48.94 47.37 47.09 44.88 50.68 49.03
DEAP-A ACC 51.48 50.99 26.48 41.08 48.94 49.73 45.90 44.09 49.91 43.86 50.05 50.59
F1 50.95 50.09 25.39 41.05 48.94 49.17 45.56 41.51 47.09 40.46 47.79 46.38
DEAP-VA ACC 24.58 27.84 16.06 17.31 25.41 23.90 24.64 20.89 25.66 19.09 30.92 20.49
F1 24.70 24.36 15.19 16.87 24.44 22.58 23.24 15.16 24.95 13.02 30.98 18.41
V-A Model Reproduction
To ensure the reliability of our benchmark, we conducted a meticulous replication of the representative EER methods using the original configurations and compared them with the results reported in the original paper. Table IV presents both our reimplementation outcomes and those reported in the original study with the gap. Additionally, we detail the specific original experimental setups, clearly illustrating the differences in experimental details among various models. All reproduced models have been integrated into our LibEER algorithm library, allowing for direct usage.
(1) The reproduced model performances are generally slightly lower than the results reported in the original paper, averaging 2.95% lower, indicating a satisfactory overall replication quality. Notably, the reproduction results for RGNN, DBN, R2G-STNN, GCBNet_BLS, and ACRNN showed more significant declines compared to the original paper. Conversely, some scenarios in MS-MDA, as well as CDCN and HSLT, exhibited improvements in replication results, while the performances of other models were comparable to those in the original study. The primary reason for these discrepancies is the lack of detailed descriptions regarding model specifications and experimental setups in some papers. We determine these parameters and settings through tuning and experience, which could introduce some biases into the results. Consequently, even when attempts are made to replicate the results as closely as possible to the descriptions in the original work, it remains challenging to achieve identical outcomes.
(2) There are significant differences in the experimental setups among various methods. The paper employs different datasets for each model, utilizes various data preprocessing techniques, executes distinct experimental tasks, and applies different data partitioning and performance evaluation methods. The substantial variability in experimental configurations prevents fair comparisons among the models. This underscores the necessity of establishing a benchmark under a unified and rational experimental framework.
V-B Model Comparison
As shown in Table IV, the implementation details of different methods vary significantly, making it difficult to fairly compare their performance and draw reliable conclusions for the advancement of the EER field. To this end, we conduct a fair comparison of ten selected methods based on the unified and reasonable benchmark proposed in Section III. The results for the two experimental tasks are reported in Tables V and VI, respectively. We select SVM [29], one of the most commonly used machine learning algorithms, as the baseline model. Its results are presented first. In the two tables, the top two methods in each scenario are highlighted using bold and underlined formatting. From these results, we make the following key observations.
Observation 1. The graph-structured information between EEG signal channels is crucial for EER tasks. Several GNN-based models achieved superior performance.
To provide a clearer illustration of the performance of each method under the benchmark, we ranked each method’s performance in each scenario. The highest-ranking method among the $n$ methods receives $n$ points, while the lowest-ranking method receives 1 point. The total score for each method across all scenarios is then accumulated, and the scoring results for both experimental tasks are presented in Fig. 7 (a) and (b).
From Fig. 7 (a), it can be observed that in the subject-dependent task, DGCNN achieved the best performance with a total score of 140. Following closely are GCBNet and its variant GCBNet_BLS, both scoring 132 points. CDCN and RGNN ranked fourth and fifth with scores of 126 and 105, respectively. In contrast, EEGNet, Tsception, and ACRNN performed relatively poorly. In the cross-subject task (see Fig. 7 (b)), the overall performance ranking of each model is similar to that in the subject-dependent scenario. Notably, the DBN model, which has a relatively simple architecture, rises to third place in the cross-subject scenario. In contrast, the MS-MDA model, specifically designed for cross-subject scenarios, shows significant variability across different datasets, resulting in a moderate ranking in the overall score.
In terms of model types, the four GNN-based methods demonstrated superior performance on both tasks, providing strong evidence that the graph-structured information between EEG signal channels is crucial for EER tasks. In contrast, the CNN-based methods exhibited varying levels of performance on both tasks. The DNN and Transformer-based methods showed average results on the subject-dependent task, while the DNN-based methods performed better on the cross-subject task. The RNN-based methods performed relatively poorly. However, for the Transformer and RNN-based methods, as only one model was selected from each category, the results may contain some occasionality. Based on the current results, we recommend that researchers in the EER field focus more on using GNN-based models when designing their models and emphasize the channel graph structure information inherent in EEG signals.
@@ V-B Model Comparison
Figure 7: The Model Performance Scoring on subject-dependent and cross-subject tasks.
(a) subject-dependent
(b) cross-subject
@@
Observation 4. The data collection methods for different EEG datasets vary, resulting in differences in data quality and task difficulty. Several methods performed better on the SEED dataset, which has a more reasonable data collection approach.
To visually illustrate the performance differences of various models across different datasets, we created radar charts for both experimental tasks, as shown in Fig. 8. From the results presented in Fig. 8, we summarize that models performed better on the three-class SEED dataset compared to the binary classification tasks in DEAP and HCI. We believe this is due to the higher label quality in SEED, where discrete emotion labels are easier for self-assessment, while continuous VAD labels tend to be more subjective. Additionally, SEED provides expectations for the type of emotions elicited by stimuli, whereas DEAP and HCI do not. However, as the number of labels increases, all models demonstrate similar performance on the four-class tasks of SEED-IV, DEAP-VA, and HCI-VA, with relatively poorer performance observed in DEAP-VA.
@@
Figure 9: Ablation study results on sample length for subject-dependent and cross-subject tasks on the SEED-IV and HCI-VA datasets.
@@
Figure 10: Ablation study results on feature type for subject-dependent and cross-subject tasks on the SEED-IV and HCI-VA datasets.
TABLE VII: Ablation study results on training methods for subject-dependent task on the SEED-IV and HCI-VA datasets.
@@
TABLE VIII: Ablation study results on training methods for cross-subject task on the SEED-IV and HCI-VA datasets.
@@ V-C Discussion of Experiment Settings
In this section, we select multiple representative methods from each kind of EER method and conduct ablation studies on three crucial factors, i.e., sample length, feature type, and evaluation methods. To be specific, SEED-IV and MAHNOB-HCI with four classification labels are taken as the dataset and SVM, DBN, HSLT, CDCN, ACRNN, and DGCNN will be selected as the representative methods. The ACC and F1 scores for both experimental scenarios are reported.
@@ V-C2 Data preprocessing Feature
In the ablation study on feature types, we utilized both DE and PSD features, along with their respective variants processed by Locality Differential Sparse (LDS), as inputs. The results are presented in Fig. 10. As illustrated in Fig. 10, in the subject-dependent scenario, the DE and PSD features demonstrate comparable performance, while both feature types processed by LDS show significant improvements. In the cross-subject scenario, the DE feature exhibited the lowest performance, whereas the other three features displayed similar efficacy. This finding underscores the utility of LDS as a valuable feature processing technique, suggesting that the selection of specific features should be adapted flexibly according to the experimental context and the unique characteristics of the model.

### v3

@@ V Experimental Results
EEGNet CDCN TSception ACRNN BiDANN R2G-STNN DGCNN GCBNet GCBNet _BLS RGNN NSAL-DGAT DBN DAN DANN MS-MDA PR-PL HSLT
SEED ACC 37.07 38.19 57.72 45.60 45.39 59.49 56.84 60.87 56.32 56.32 59.45 64.40 36.16 44.49 63.89 64.00 56.33 56.00
(15.54) (17.25) (16.50) (8.71) (10.09) (15.92) (19.20) (15.62) (14.73) (9.91) (15.28) (19.00) (6.32) (7.68) (1.37) (18.75) (4.69) (14.76)
F1 33.45 31.83 58.66 43.54 42.37 48.18 51.61 57.22 55.12 51.43 56.31 58.47 22.67 33.79 53.37 57.35 48.02 55.75
(10.98) (12.33) (13.47) (9.72) (9.74) (18.23) (13.40) (16.41) (12.88) (6.54) (12.56) (19.50) (2.57) (5.60) (2.98) (18.67) (2.36) (14.73)
SEED-IV ACC 28.98 28.19 31.03 34.19 31.97 42.28 42.39 42.54 32.27 40.54 48.23 52.29 36.82 38.24 25.55 56.07 43.42 30.33
(4.09) (3.82) (9.45) (6.99) (7.02) (15.67) (16.02) (22.09) (11.43) (7.33) (16.14) (16.89) (12.81) (11.86) (2.54) (15.45) (5.48) (5.47)
F1 24.56 28.35 27.01 26.83 18.82 37.31 42.76 43.10 32.89 42.73 45.43 45.39 32.60 26.04 11.99 48.68 42.37 11.64
(5.31) (5.44) (6.63) (6.84) (6.80) (14.14) (17.54) (23.71) (12.53) (5.89) (14.89) (13.29) (12.81) (11.74) (1.96) (13.78) (2.39) (2.34)
HCI-V ACC 70.33 57.06 67.69 57.36 54.53 73.92 64.76 63.19 65.16 71.06 64.94 64.04 69.27 61.50 63.61 67.64 68.79 66.94
(8.74) (8.14) (8.15) (10.99) (9.87) (10.56) (13.28) (8.13) (6.76) (13.33) (6.72) (6.67) (8.09) (11.49) (6.57) (7.88) (10.43) (5.52)
F1 66.36 53.83 62.67 54.76 52.58 66.02 59.54 58.75 63.32 61.83 52.50 38.96 65.51 50.97 38.80 53.70 54.90 62.22
(5.89) (5.59) (10.57) (9.28) (4.22) (14.04) (13.41) (10.52) (2.33) (9.43) (12.87) (2.26) (12.54) (9.86) (2.22) (3.41) (14.21) (4.53)
HCI-A ACC 54.41 54.70 55.93 52.30 51.23 49.98 56.21 59.42 52.21 60.85 57.36 51.00 57.50 56.42 49.06 57.23 55.48 49.48
(14.26) (14.46) (19.96) (16.16) (16.66) (22.61) (13.33) (19.93) (10.83) (16.25) (22.72) (23.13) (9.23) (18.05) (23.43) (21.98) (16.73) (18.17)
F1 52.91 54.02 55.15 50.25 49.45 48.53 55.24 57.02 51.84 56.26 56.90 32.41 56.24 45.35 34.68 55.68 54.45 47.43
(11.20) (10.21) (11.60) (11.56) (13.70) (12.59) (10.27) (11.56) (12.32) (12.93) (11.78) (10.08) (5.51) (12.41) (10.36) (9.65) (9.92) (13.90)
HCI-VA ACC 31.56 34.84 30.09 26.99 27.21 36.24 38.25 41.54 43.02 36.92 42.63 38.66 28.30 40.35 28.00 44.92 44.24 35.07
(9.12) (9.32) (9.81) (8.81) (12.14) (19.56) (11.10) (9.82) (12.60) (18.54) (17.44) (22.77) (19.84) (22.83) (18.26) (20.67) (17.24) (13.21)
F1 26.29 27.96 24.71 21.95 20.92 24.42 28.48 38.08 36.07 32.99 25.28 22.13 27.58 20.71 13.12 21.65 22.93 27.19
(3.57) (3.12) (2.47) (6.24) (4.61) (5.91) (5.52) (2.43) (7.58) (12.41) (6.22) (8.92) (8.97) (7.13) (6.28) (8.94) (2.29) (5.64)
DEAP-V ACC 49.58 52.36 57.78 54.44 51.94 53.49 49.97 49.91 53.58 52.34 52.65 52.22 55.02 51.88 44.32 54.82 51.77 56.56
(6.77) (6.81) (5.00) (6.33) (7.44) (7.07) (7.19) (7.48) (8.72) (4.51) (8.09) (11.83) (7.19) (7.90) (5.94) (10.42) (10.60) (5.79)
F1 48.49 49.74 57.72 48.94 47.37 45.96 47.75 47.09 50.68 49.03 50.22 45.10 53.14 44.61 30.61 50.60 40.21 56.56
(6.34) (6.32) (6.00) (7.27) (6.66) (9.86) (7.75) (7.31) (11.20) (5.28) (6.23) (10.30) (7.16) (9.50) (2.58) (9.65) (13.48) (2.21)
DEAP-A ACC 51.48 48.94 49.73 45.90 44.09 50.07 49.13 49.91 50.05 50.59 46.85 50.94 50.99 36.86 49.32 26.48 45.01 41.08
(7.76) (7.72) (9.38) (8.16) (16.08) (9.01) (7.49) (9.27) (10.60) (19.22) (16.25) (9.60) (13.86) (10.64) (16.56) (13.21) (13.05) (14.32)
F1 50.95 48.94 49.17 45.56 41.51 46.87 47.18 47.09 47.79 46.38 43.90 50.75 50.09 32.68 33.26 25.39 37.26 41.05
(6.88) (6.75) (9.73) (7.29) (7.69) (8.20) (5.83) (6.12) (6.37) (11.85) (10.58) (8.46) (12.74) (9.55) (6.78) (10.91) (10.96) (13.94)
DEAP-VA ACC 24.58 25.41 23.90 24.64 20.89 29.30 28.02 25.66 30.92 20.49 22.40 26.44 27.84 23.87 26.44 16.06 22.85 17.31
(8.90) (8.86) (9.11) (8.69) (14.62) (8.34) (9.12) (8.22) (12.73) (11.13) (12.36) (14.74) (13.66) (6.08) (14.74) (14.23) (8.14) (14.77)
F1 24.70 24.44 22.58 23.24 15.16 23.36 23.14 24.95 30.98 18.41 16.85 9.97 24.36 22.33 9.97 15.19 15.49 16.87
(6.98) (6.49) (6.88) (7.46) (7.56) (4.24) (8.70) (5.32) (5.27) (10.33) (6.13) (4.58) (5.53) (5.47) (4.58) (5.02) (6.62) (5.17)
SEED-V ACC 28.03 32.27 33.84 32.63 27.03 36.58 37.23 36.25 30.84 35.68 36.45 41.56 32.18 20.82 23.99 44.33 33.77 30.59
(10.91) (3.57) (9.87) (4.90) (12.76) (12.89) (12.47) (2.61) (6.51) (13.22) (5.65) (13.19) (6.56) (5.45) (3.98) (12.09) (2.24) (12.13)
F1 19.54 22.22 27.45 25.33 18.27 32.72 31.48 29.77 27.70 30.02 31.93 35.33 25.29 14.97 17.96 38.28 29.64 26.54
(11.93) (3.42) (11.00) (5.61) (11.65) (12.71) (13.62) (2.25) (3.64) (16.43) (10.42) (12.84) (7.71) (4.54) (4.08) (13.48) (3.00) (5.87)
MPED ACC 19.56 20.33 17.83 17.65 17.53 25.75 22.29 19.64 18.82 19.01 18.85 19.07 17.83 19.10 17.77 20.67 18.76 17.94
(3.29) (5.44) (1.09) (1.14) (2.41) (7.97) (5.40) (5.46) (2.69) (1.63) (3.35) (0.57) (1.09) (1.89) (1.44) (2.15) (0.39) (4.21)
F1 16.36 17.27 15.31 15.65 15.28 21.26 18.19 16.25 15.42 15.42 15.22 15.55 15.31 17.17 15.54 16.79 16.10 15.27
(3.75) (3.60) (1.19) (1.70) (4.11) (6.81) (5.44) (5.56) (2.89) (2.07) (2.37) (1.48) (1.19) (1.89) (1.50) (3.51) (0.60) (3.92)
@@
Figure 5: The model performance scoring on subject-dependent and cross-subject tasks based on ACC and F1.
V-A Model Reproduction
To ensure the reliability of our benchmark, we conducted a systematic reproduction of the representative EER methods by strictly following the experimental settings reported in their original papers, including model configurations and dataset usage. Table IV presents both our reproduction results and the original reported results, along with the performance gap. Considering that SEED and DEAP are the two most widely used public datasets, and all included methods have been evaluated on at least one of them, we reproduced the results specifically on SEED and DEAP. If the gap was within 10%, we regarded the reproduced implementation as sufficiently reliable for subsequent unified benchmark comparisons. In addition, Table IV provides the specific original experimental setups used in the reproduction experiments, clarifying the differences in settings across different models compared to our unified benchmark. All verified models have been integrated into our LibEER algorithm library for direct use.
(1) The reproduced model performances are generally slightly lower than the results reported in the original paper, averaging 3.85% lower, with a maximum difference of less than 10%, indicating a satisfactory overall replication quality. The reproduction results for ACRNN, R2G-STNN, GCBNet_BLS, RGNN, DBN, and DAN showed more significant declines compared to the original papers ($>$5%). The reproduction results for BiDANN, DGCNN, GCBNet, NSAL-DGAT, and PR-PL showed smaller deviations ($<$5%). Conversely, some scenarios in CDCN, TSception, MS-MDA, and HSLT exhibited improvements in the reproduced results. The primary reason for these discrepancies is the lack of detailed descriptions regarding model specifications and experimental setups in some papers. We determine these parameters and settings through tuning and experience, which could introduce some biases into the results. Consequently, even when attempts are made to replicate the results as closely as possible to the descriptions in the original work, it remains challenging to achieve identical outcomes.
(2) There are substantial inconsistencies in the experimental setups reported in the original papers, including differences in datasets, preprocessing, tasks, and evaluation protocols. Notably, all original settings do not adopt a three-way split (train/validation/test), which increases the risk of overfitting and inflating performance due to the lack of validation-based model selection. This substantial variability in experimental configurations prevents fair comparisons among the models. Therefore, we established a unified benchmark under a consistent and rational experimental framework, as described in Section III, where all methods are evaluated using standardized preprocessing, dataset usage, a three-way split, and consistent evaluation metrics.
V-B Model Comparison
As shown in Table IV, the implementation details of different methods vary significantly, making it difficult to fairly compare their performance and draw reliable conclusions for the advancement of the EER field. To this end, we conduct a fair comparison of seventeen selected methods based on the unified and reasonable benchmark proposed in Section III. The results for the two experimental tasks are reported in Tables V and VI, which clearly present results that differ significantly from those in Table IV. We select SVM [36], one of the most commonly used machine learning algorithms, as the baseline model. Its results are presented first. In the two tables, the top two methods in each scenario are highlighted using bold and underlined formatting. From these results, we make the following key observations.
Observation 1. Modeling the spatiotemporal information of EEG signals is crucial for EER tasks. Several GNN-based and RNN-based models achieved superior performance. In addition, integrating transfer learning can effectively alleviate inter-subject variability in EEG data, demonstrating considerable potential.
@@ V-B Model Comparison
To provide a clearer illustration of the performance of each method under the benchmark, we designed a scoring scheme that simultaneously considers accuracy and F1-score across all datasets. Specifically, for ten datasets (SEED, SEED-IV, HCI-V, HCI-A, HCI-VA, DEAP-V, DEAP-A, DEAP-VA, SEED-V, MPED), all $n$ methods are evaluated based on both accuracy and F1 score, yielding a total of 20 rankings (10 datasets × 2 metrics). In each ranking, the top-performing method receives $n$ points, while the lowest receives 1 point. The final score of each method is computed as the sum of its points across all 20 rankings, as shown in Fig. 5 (a) and (b).
From Fig. 5 (a), it can be observed that in the subject-dependent task, DGCNN achieved the best performance with a total score of 275. Following closely are GCBNet and its variant GCBNet_BLS, scoring 259 and 255 points, respectively. CDCN and BiDANN ranked fourth and fifth with scores of 235 and 234, respectively. In contrast, EEGNet, Tsception, ACRNN, DAN, and DANN performed relatively poorly. In the cross-subject task (see Fig. 5 (b)), the overall performance ranking of each model is similar to that in the subject-dependent scenario. Several GNN-based methods still achieved relatively superior performance. BiDANN and R2G-STNN, two models based on RNN and transfer learning that already performed well in subject-dependent settings, achieved even better performance in the cross-subject scenario, ranking second and third, respectively. EEGNet and DBN obtained better results on this task compared to their performance in subject-dependent settings. The MS-MDA model, specifically designed for cross-subject scenarios, resulted in a moderate ranking in the overall score.
In terms of model types, the four GNN-based methods demonstrated superior performance on both tasks, providing strong evidence that the graph-structured information between EEG signal channels is crucial for EER tasks. In contrast, the CNN-based methods exhibited varying levels of performance on both tasks. The DNN and Transformer-based methods showed relatively weaker overall performance, though some individual models demonstrated promising results on specific tasks. However, for the Transformer-based methods, as only one model was selected, the results may contain some occasionality. RNN-based models exhibited notable performance divergence, with both BiDANN and R2G-STNN ranking within the top five and even top three across both tasks, while ACRNN consistently ranked last. This may be attributed to the incorporation of transfer learning in BiDANN and R2G-STNN, which effectively models inter-sample variability. However, the two DNN-based transfer learning models, DAN and DANN, performed poorly, indicating that beyond employing transfer learning, it is equally important to capture the intrinsic spatiotemporal characteristics of EEG signals through architectures such as RNNs and GNNs. Based on the current results, we recommend that researchers in the EER field prioritize models capable of capturing the spatiotemporal dynamics of EEG signals, such as those based on RNN or GNN, and consider incorporating transfer learning techniques to enhance model generalizability across subjects.
@@
Figure 7: Experimental results on sample length for subject-dependent and cross-subject tasks on the SEED-IV and HCI-VA datasets.
@@
Figure 8: Experimental results on feature type for subject-dependent and cross-subject tasks on the SEED-IV and HCI-VA datasets.
TABLE VII: Experimental results on training methods for subject-dependent task on the SEED-IV and HCI-VA datasets.
@@
TABLE VIII: Experimental results on training methods for cross-subject task on the SEED-IV and HCI-VA datasets.
@@
Observation 4. The data collection methods for different EEG datasets vary, resulting in differences in data quality and task difficulty. Several methods performed better on the SEED series datasets, which feature a more reasonable data collection approach.
To visually illustrate the performance differences of various models across different datasets, we created radar charts for both experimental tasks with ACC and F1 score, as shown in Fig. 6. From the results, we summarize that models performed better on the three-class SEED dataset compared to the binary classification tasks in DEAP and HCI. We believe this is due to the higher label quality in SEED, where discrete emotion labels are easier for self-assessment, while continuous Valence-Arousal-Dominance labels tend to be more subjective. Additionally, SEED provides expectations for the type of emotions elicited by stimuli, whereas DEAP and HCI do not. As the number of labels increases, various models achieve relatively similar average accuracy on SEED-IV, HCI-VA, and SEED-V, with only DEAP-VA showing notably lower performance. However, the performance differences among models are more pronounced on SEED-IV and SEED-V compared to the other two datasets. Moreover, due to the more balanced label distribution in the SEED series datasets, their F1 scores are significantly higher than those of the other datasets. The MPED dataset, which contains the largest number of emotion categories, poses significant challenges to most models, indicating that fine-grained recognition of human emotions in the EER field still requires further investigation.
(a) Subject-dependent task on the SEED-IV dataset
(b) Cross-subject task on the SEED-IV dataset
(c) Subject-dependent task on the HCI-VA dataset
(d) Cross-subject task on the HCI-VA dataset
Figure 9: Experimental results on random seed for subject-dependent and cross-subject tasks on the SEED-IV and HCI-VA datasets.
@@ V-C Discussion of Experiment Settings
In this section, we select multiple representative methods from each kind of EER method and conduct comparative experiments on four crucial factors, i.e., sample length, feature type, evaluation methods and random seeds. To be specific, SEED-IV and MAHNOB-HCI with four classification labels are taken as the dataset and SVM, DBN, HSLT, CDCN, ACRNN, and DGCNN will be selected as the representative methods. The ACC and F1 scores for both experimental scenarios are reported.
@@ V-C2 Data preprocessing Feature
In the comparative experiments on feature types, we utilized both DE and PSD features, along with their respective variants processed by Locality Differential Sparse (LDS), as inputs. The results are presented in Fig. 8. As illustrated in Fig. 8, in the subject-dependent scenario, the DE and PSD features demonstrate comparable performance, while both feature types processed by LDS show significant improvements. In the cross-subject scenario, the DE feature exhibited the lowest performance, whereas the other three features displayed similar efficacy. This finding highlights LDS as a valuable feature processing technique, suggesting that the selection of specific features should be adapted flexibly according to the experimental context and the unique characteristics of the model.